Skateboard Corp. has 7 million shares of common stock outstanding. The current share price is $79, and the book value per share is $6. The company also has two bond issues outstanding. The first bond issue has a face value of $70 million, has a 8 percent coupon, and sells for 94 percent of par. The second issue has a face value of $40 million, has a 9 percent coupon, and sells for 107 percent of par. The first issue matures in 23 years, the second in 6 years.

   

The most recent dividend was $5.1 and the dividend growth rate is 9 percent. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 34 percent.

   

What is the company's WACC?
Answer #1

Price = 01/(ke-g) 5.1*1.09 D1 = Ke-g= Ke = $79.00 9.00% $5.56 0.0704 16.04% 0.1604 Cost of equity = After tax cost of First i
Answer #2

Correction: in your first excel chart, column "weight x YTM" should read "weight x After tax%"  Because YTM would be a different % ... YTM =PRE TAX

Did you check to make sure your final answer was correct? thank

answered by: Kyle Hardwick